OpenStudy (anonymous):

Solve the system of equations. y= x^2/(x-1) +1 y= 1/(x-1) someone help, thanks!

OpenStudy (anonymous):

set y1 = y2 and solve for x first

hero (hero):

@TheRealMeeeee , @turnupcity probably needs help with it because the system posted includes fractions.

OpenStudy (anonymous):

all i know was to set y1 = y2 and solve for x but don't really know how to do it

hero (hero):

After setting \(y_1 = y_2\) you have: \[\frac{1}{x - 1} = \frac{x^2}{x - 1} + 1\] Right @turnupcity ?

OpenStudy (anonymous):

yeah thats exactly it

hero (hero):

Okay, now since we have two fractions with the same denominator we can add or subtract them right?

OpenStudy (anonymous):

yes..

hero (hero):

Okay, well, the next step before we do that is to subtract 1/(x - 1) from both sides: \[0 = \frac{x^2}{x - 1} - \frac{1}{x - 1} + 1\]

hero (hero):

Hopefully you understood how I got that.

OpenStudy (anonymous):

yes so far i do

hero (hero):

Because the next step would be to combine the fractions with like denominators: \[0 = \frac{x^2 - 1}{x - 1} + 1\]

hero (hero):

At this point you should notice that the numerator has the term \(x^2 - 1\) which is a difference of squares. In other words, we can factor it to get \((x + 1)(x - 1)\)

OpenStudy (anonymous):

okay.. lol how would do that

hero (hero):

Keep in mind that in general, if you have a difference of squares \(a^2 - b^2\) it factors to \((a + b)(a - b)\)

hero (hero):

Because in general \(a^2 - b^2 = (a + b)(a - b)\)

hero (hero):

So what we have is \[0 = \frac{(x + 1)(x - 1)}{x - 1} + 1\]

OpenStudy (anonymous):

so would we multiply (x+1) on the bottom?

OpenStudy (anonymous):

or add (x+1)?

hero (hero):

No, you would not. At this point you should see that you have the same number on the top and bottom \((x - 1)\)

OpenStudy (anonymous):

so it would cancel out?

hero (hero):

Exactly

OpenStudy (anonymous):

so it would leave it at y=(x+1)+1?

hero (hero):

No, if you look at my last step you'll notice that the left side was zero. Why did you replace zero with y?

OpenStudy (anonymous):

oops didnt see that, so does x =-2?

hero (hero):

Exactly. Now if you plug x back in to the original system, you'll be able to find the value of y

OpenStudy (anonymous):

thank you, apperciate it!

hero (hero):

Let me know what you get for y

hero (hero):

If you run into any trouble with it, let me know. I can easily and quickly assist.

OpenStudy (anonymous):

i got -.33333333? i might have typed it wrong in my calc

hero (hero):

I was hoping you would do it without the need of a calc.

hero (hero):

To do it without a calc, here's what you would do. First realize that it would be easier to solve for y if we use this equation: \[y = \frac{1}{x - 1}\] So plug in x = -2... \[y = \frac{1}{-2 - 1}\] \[y = \frac{1}{-3}\] \[y = -\frac{1}{3}\]

hero (hero):

Basically, for finding y, you shouldn't need a calc if all you're doing is plugging back in.

OpenStudy (anonymous):

oh that makes sense, yeah i guess i didnt my calc for that

hero (hero):

So the solution point is \[(x,y) = \left(-2, -\frac{1}{3}\right)\]